This is pretty easy, but I am drawing a blank:
How do you divide 962 by 12?
(I am unsure how to handle the remainder.)
If I use the calculator, the answer is 80.1, but if I do it on paper, I cannot get the answer to correspond.
Thanks...

Answer by Alan3354(69443)

962/12 = 80 with a remainder of 2
Or, 962/12 = 80 1/6
Or, 962/12 = 80.1666666...